A random variable Q has pdf ?q(q) = {\frac{1}{4}|x| − 2 ≤ Q ≤ 2; 0 else} .

(1) Find ?[Q], and Var(Q).

(2) Let Z = Q2. Find ?z(Z).

(3) Find P{Q > 1}

(4) Find conditional PDF of Q given B= {Q > 1} i.e., find ?Q|?(Q).
Answer #1

2 = 8 - Var (8) = E(87) - (E(81) 2- or 2. = P(ruz F (2) P(2.42) - Plan (2) - P 1815 da) Lo Lda fa (a) da uz Vz 2 VZ -adat Sad(0) 4)4318 8(9 十(0(372 Pacocs @>) P103) Pi<952) P1K022) )) 1